OpenStudy (anonymous):

Find the product. (Multiplying Rational Expressions) v - 12 over 10 times - v - 2 over 4 Help/an explanation would be greatly appreciated. :)

OpenStudy (anonymous):

All you have to do is multiply the numerators and the denominators

OpenStudy (anonymous):

I did that, but I got a completely different answer from the book, which has an answer key. I'm pretty sure I have to factor something, but I'm not sure. The book says the answer is \[-(v-3) \over 5(v+1)\].

OpenStudy (anonymous):

can you show your work and what you got?

OpenStudy (anonymous):

The drawing feature isn't being cooperative, but I'll try to explain it as best I can. I multiplied (v-12) and -v-2 and got v+24 over 40. And I'm somewhat stuck there right now.

OpenStudy (anonymous):

i'm doing it out on on paper and i got (-v^2+10v+24)/40 so the answer in your bok is stumping me

OpenStudy (anonymous):

Hm.. I'll keep trying. Maybe the answer was written incorrectly. Thanks for the assistance, though!

OpenStudy (anonymous):

no problem. I'd suggest bumping this and see if anyone else can figure it out.

OpenStudy (pfenn1):

@ChukRock , I got the same thing you did. @Hologram are you sure that you copied the problem correctly from the book and that you are looking at the correct answer?

OpenStudy (anonymous):

My question was a simplified version of the original problem, so it's likely that I simplified the problem incorrectly. Would it be better to post the original instead? Sorry for the confusion; this is a totally new concept.

OpenStudy (pfenn1):

Yes, try to post the original

OpenStudy (anonymous):

Alright, the original is \[v ^{2}+v-12 \over 5v+1 \] times \[-v-2 \over v ^{2}+5v +4\].

OpenStudy (pfenn1):

Is the denominator of that first term 5v+1 or is it v+1?

OpenStudy (anonymous):

Sorry, it's 5v+10. Missed that.

OpenStudy (pfenn1):

Ok, factor\[v ^{2}+v-12=\]and\[v ^{2}+5v+4=\] and \[5v+10=\] and plug back into the original equation. I think you will then get the answer from the book.

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (pfenn1):

you're welcome